Can anyone help? I need to solve this problem with work shown:
Show a complete solution to each problem.
Big Plane. A six passenger plane cruises at 180 mph in calm air. If the plane flies 7 miles with the wind in the same amount of time as it flies 5 miles against the wind then what is the wind speed?

Answer by stanbon(75887)

A six passenger plane cruises at 180 mph in calm air. If the plane flies 7 miles with the wind in the same amount of time as it flies 5 miles against the wind then what is the wind speed?
With the wind DATA:
distance=7 mi; rate=180+wind; time=d/r=7/(180+w)
Against the wind DATA:
distance=5 mi; rate=180-wind; time=d/r=7/(180-w)
EQUATION:
time with = time against
7/(180+w) = 5/(180)-w
7(180)-w =5(180+w)
1260-7w=900+5w
12w=360
w=30 mph (speed of wind)
Cheers,
Stan H.